# SH2: Self-Highlighted Hesitation Helps You Decode More Truthfully

### 11 Jan 2024 v1

@@ Abstract
Large language models (LLMs) demonstrate great performance in text generation. However, LLMs are still suffering from hallucinations. In this work, we propose an inference-time method, Self-Highlighted Hesitation (SH2), to help LLMs decode more truthfully. SH2 is based on a simple fact rooted in information theory that for an LLM, the tokens predicted with lower probabilities are prone to be more informative than others. Our analysis shows that the tokens assigned with lower probabilities by an LLM are more likely to be closely related to factual information, such as nouns, proper nouns, and adjectives. Therefore, we propose to “highlight” the factual information by selecting the tokens with the lowest probabilities and concatenating them to the original context, thus forcing the model to repeatedly read and hesitate on these tokens before generation. During decoding, we also adopt contrastive decoding to emphasize the difference in the output probabilities brought by the hesitation. Experimental results demonstrate that our SH2, requiring no additional data or models, can effectively help LLMs elicit factual knowledge and distinguish hallucinated contexts. Significant and consistent improvements are achieved by SH2 for LLaMA-7b and LLaMA2-7b on multiple hallucination tasks.  [^1] We will release our code for reproducibility later.
@@ 1 Introduction
Depending on massive training corpora, large language models (LLMs) have made significant progress in natural language understanding and text generation (Touvron et al., 2023a, Sun et al., 2023, OpenAI, 2023). However, during reasoning and generation, LLMs could suffer from hallucinations and generate non-factual answers (Zhang et al., 2023b).
@@ 2 Related Work
The idea of using extra decoding steps while predicting harder tokens can be dated back to Adaptive Computation Time proposed by Graves (2017). Shin et al. (2020) introduce AUTOPROMPT to combine the original inputs with trigger tokens to elicit knowledge from pretrained models. Wiegreffe et al. (2022) investigate that additional steps to generate rationales could lead to a more faithful model. More recently, Goyal et al. (2023) have also demonstrated that for LLMs, inserting extra computation steps to allow the model to pause before generation can improve the performance on question answering and reasoning tasks. Nevertheless, their method only works when the model is both pre-trained and finetuned with pauses.
@@ 3.2 Relation between Key Tokens and Factual Information
It can be observed from Figure 2 that content words such as adjectives (JJ), nouns (NN), proper nouns (NNP), adverbs (RB) and conjugated verbs (VBD, VBG, VBP) are more frequent in the hardest words we extract, as the light color in the heatmap indicates. On the other hand, function words such as conjunctions (CC), determiners (DT), and prepositions (IN) are less frequent, as indicated by the darker colors. [^2] The base form of the verbs (VB) seems to be an exception. It is probably because VBs have less information load than conjugated verb forms, as the latter also encode tense, aspect, and other information about the verb. This suggests that the normalized top-$\eta$ recall is positively correlated with factual informativeness. Therefore, picking out key tokens that carry more content, rather than those that serve as grammatical functions, could intuitively help the model to focus on factual information. Our method can thus be understood as a way to “highlight” the content words to LLMs, potentially making them less likely to hallucinate, the effect of which will be evaluated in our experiments. More cases will be given in Appendix C.2.
@@ 3.3 Construction of Hesitations
First, construct a candidate key-token set by selecting $\eta\cdot n$ tokens with the lowest prediction probabilities of $\hat{P}(X)$, where $\eta\in(0,1)$ is the sampling proportion we preset for $X$. Then, derive the target key-token set $T(X)$ by randomly drop out $\lambda\cdot\eta\cdot n$ tokens from the candidate set, where $\lambda\in[0,1]$ is the drop out rate.
For long documents, to avoid $T(X)$ got dominated by one or two POS tags, we introduce the drop out rate $\lambda$ to randomly select key tokens among a larger pool of candidates, whose size is determined by $\eta$. As a result, $(1-\lambda)\cdot\eta\cdot n$ tokens are retained in $T(X)$ to fill as input during the hesitation steps.
@@ 4.2 Experimental Settings
We apply our SH2 on LLaMA-7b (Touvron et al., 2023a) and LLaMA2-7b (Touvron et al., 2023b). We compare it with other SOTA (state-of-the-art) methods, including Alpaca (Taori et al., 2023), ITI (Li et al., 2023b), 13b-CD (Chuang et al., 2023) and DoLa (Chuang et al., 2023). All of these baselines use LLaMA-7b as their backbone. It should be noted that ITI uses 5% samples of TruthfulQA in few-shot prompting. Additionally, LLaMA-13b is used in 13b-CD to be contrasted with the 7b model. We implement Alpaca and DoLa following the official instructions and report the evaluation results of our implementation.
@@
Since there are only about a dozen tokens in each question of TruthfulQA, the sample proportion $\eta$ of LLaMA-7b is set to be 10% and 40% for the discrimination track and the generation track respectively. And the drop rate $\lambda$ is set to be 0. While for HaluEval-Sum, which has about a thousand tokens in each document, $\eta$ and $\lambda$ are set to be 6% and 0.33 for LLaMA-7b. The settings of the hyperparameters $\eta$, $\lambda$ and $\alpha$ for LLaMA-7b and LLaMA2-7b on each task are summarized in Appendix A.
@@ 4.3 Main Results
The results on the three benchmarks are shown in Table 1. Our SH2 obtained significant and consistent improvements for both LLaMA-7b and LLaMA2-7b. When compared with other SFT or decoding reformulation methods, our method outperforms them in most of the metrics of the benchmarks. There is no additional data or model used in our SH2. We ask models to select the hardest tokens and hesitate on them. Although LLaMA2 has been aligned to generate truthful answers during its training procedure, it still benefits significantly from our inference-time method.
@@
(a) Average scores (%) for different highlighted tokens on the sampling proportion $\eta$ ranging from 1% to 8%. The errorbar denotes standard deviations.
(b) Harmonic accuracy (%) for different highlighted tokens with respect to the sampling proportion $\eta$.
(c) Precision, recall and F1 scores (%) for different highlighted tokens. The scores of the vanilla LLaMA2-7b are obtained by evaluating on the whole dataset of HaluEval-Sum.
Figure 3: Different choices of highlighted tokens by LLaMA2-7b.
@@ 4.4 Bias in HaluEval-Sum
The great gap between Acc-A and Acc-H denotes LLMs’ bias in hallucinated summaries and right summaries even for LLaMA2-7b. Our findings suggest that hallucination could not be solved fundamentally through training.
@@
Nevertheless, our method successfully mitigates this imbalance and benefits the general performance. It can improve the discriminant ability of the model comprehensively.
@@ 5.1 Choices of Highlighted tokens
(a) Different numbers of highlighted tokens.
(b) Different numbers of pause words.
Figure 4: Effect of contrastive decoding on hesitations. The dashed line in gray represents the MC2 score of the vanilla LLaMA-7b. Dashed lines in other colors represent the MC2 scores for the standard contrastive decoding with different numbers of highlighted tokens (Figure 4a) or pause words (Figure 4b).
In contrast to key tokens that are the hardest to predict, we sample the easiest tokens with the highest prediction probability. Additionally, we also randomly sample the same number of tokens for comparison. We conduct experiments with LLaMA2-7b on 1000 samples of HaluEval-Sum. The sampling proportion $\eta$ ranges from 1% to 8% with the step of %1. Average scores and standard deviations are calculated for the three choices of highlighted tokens.
Moreover, we compare precision, recall and F1 scores for the three choices of highlighted tokens in Figure 3c. The disturbance brought by highlighting the easiest tokens or random tokens hurts the precision of LLMs, while the hardest tokens are beneficial. It can also be concluded from the figure that the manner of hesitations helps mitigate the bias in HaluEval-Sum.
5.2 Manners of Hesitations
In addition to underlining key tokens in the input text, we can also repeat the text or pause. Experiments are conducted on TruthfulQA with LLaMA-7b. For the repetition manner, we can repeat the questions in hesitations. While for the pausing manner, key tokens $T_{2}(X)$ are replaced by several pause words (".") in hesitations. The performance of pausing hesitations is evaluated with 3, 6, 9 and 12 pause words. MC2 scores for different numbers of pause words are shown in Figure 4b. 6 pause words are used in hesitations in the following study.
@@ 5.3 Effect of Contrastive Decoding
Our method does not use the original form of contrastive decoding. The difference of confidence brought by hesitations, which is the item of contrastive decoding, is traded off with the base probability. We conduct experiments on TruthfulQA with different values of the adjustment hyperparameter $\alpha$ to study the effect of contrastive decoding (CD).
The effect of contrastive decoding with different numbers of highlighted tokens is shown in Figure 4a. Since there are only about a dozen tokens in each question of TruthfulQA, the hesitation contains only one token when sampling the top 10% hardest tokens. LLaMA-7b benefits a lot from the one or two highlighted tokens.
When $\alpha$ is set to be 0, the model directly decodes with the input text followed by the hesitation. The MC2 score even drops a little compared to the vanilla LLaMA-7b. As more weight is put on the difference brought by hesitations, the model obtains significant improvement. Specifically, when the weight is big enough to ignore the base probability, the decoding procedure is equivalent to standard contrastive decoding without the base probability as the equation 5. It can be concluded from the figure that, the highlighted tokens make a difference in the confidence to identify factual knowledge.
As shown in Figure 4b, we also studied the effect of contrastive decoding with different numbers of pause words. A similar conclusion can be observed from the figure.
A case study is given in Figure 5 to illustrate the effect of contrastive decoding. We calculate the probability by standard decoding without and with the hesitation for each option, and derive the difference in probability. It demonstrates that CD helps LLMs distinguish correct answers from false ones. It plays a role in separating the positive and negative answers.
@@ 6 Conclusion
In this paper, we study the bias in capturing important factual information for large language models and introduce a novel inference-time method SH2. Our method proposes to highlight the key tokens that are hard for LLMs to predict and construct hesitations with these informative tokens. To help LLMs capture critical semantic information, the decoding procedure is reformulated with the probability difference brought by hesitations. Comprehensive experiments and analysis are conducted on multiple tasks. SH2 significantly improves the truthfulness of LLMs with no need of external data or models. In the age of LLMs, our SH2 could inspire a new paradigm for faithful LLMs.
Appendix A Hyperparameter Selection
The settings of the hyperparameters $\eta$, $\lambda$ and $\alpha$ for LLaMA-7b Touvron et al. (2023a) and LLaMA2-7b Touvron et al. (2023b) on each task are summarized in Table 4.
@@ Appendix B Multiple-Choice Metrics
Appendix C Key Tokens
C.1 Normalized Top-$\eta$ Recall
We also calculated the normalized top-$\eta$ recall between the hardest part and the whole document for the top 20 most frequent POS tags by LLaMA-7b through the equation (3). The heatmap is shown in Figure 6. The same statistical laws can be concluded as those of LLaMA2-7b in Section 3.2.
Figure 6: The normalized top-$\eta$ recall for different POS tags by LLaMA-7b.
C.2 Visualization of Token Probabilities
@@
Figure 7: Visualization of token probabilities estimated by LLaMA-7b and LLaMA2-7b on case 1. The color coding in the figure represents the token generation probabilities. Tokens with lower probabilities are colored more red, while tokens with higher probabilities are colored more green.
@@
Figure 8: Visualization of token probabilities estimated by LLaMA-7b and LLaMA2-7b on case 2. The color coding in the figure represents the token generation probabilities. A spectrum from red to green is utilized to represent a range from lower to higher generation probabilities, respectively.
@@
Figure 9: Visualization of token probabilities estimated by LLaMA-7b and LLaMA2-7b on case 3. The color coding in the figure represents the token generation probabilities. Tokens with lower probabilities are colored more red, while tokens with higher probabilities are colored more green.
Figure 7 to Figure 9 illustrate three text examples where the background color of each token denotes its generation probability. A spectrum from red to green is utilized to represent probabilities ranging from low to high.
These examples reveal a tendency for lower probability tokens to predominantly consist of content words. However, instances exist where determiners and prepositions also exhibit lower probabilities, as exemplified in Figure 7a with the words “all” and “on” in line 5, and “in” in line 14. This can be attributed to the interchangeable use of different words to convey identical semantic content at the beginning of a sentence or sub-clause.
Besides, tokens with higher probabilities are typically from function words, although exceptions are noted with certain content words demonstrating high probabilities due to the extensive knowledge memorized by LLMs, such as “typhus” and “15” in Figure 9a. Additionally, some non-initial tokens such as “nesday” and “CC” in lines 1 and 3 of Figure 8a, exhibit high generation probabilities. It reflects the determinative role of initial tokens in setting the context for subsequent non-initial tokens.
Comparing the prediction probabilities of LLaMA-7b and LLaMA2-7b, we did not observe significant difference. Although LLaMA2-7b was trained with 40% more tokens compared to LLaMA-7b, there is still considerable room for improvement in its grasp of factual knowledge. Therefore, it is beneficial to incorporate hesitations to make the model pay more attention to these informative key tokens.

### 21 Feb 2024 v2

@@ Abstract
Large language models (LLMs) demonstrate great performance in text generation. However, LLMs are still suffering from hallucinations. In this work, we propose an inference-time method, Self-Highlighted Hesitation (SH2), to help LLMs decode more truthfully. SH2 is based on a simple fact rooted in information theory that for an LLM, the tokens predicted with lower probabilities are prone to be more informative than others. Our analysis shows that the tokens assigned with lower probabilities by an LLM are more likely to be closely related to factual information, such as nouns, proper nouns, and adjectives. Therefore, we propose to “highlight” the factual information by selecting the tokens with the lowest probabilities and concatenating them to the original context, thus forcing the model to repeatedly read and hesitate on these tokens before generation. During decoding, we also adopt contrastive decoding to emphasize the difference in the output probabilities brought by the hesitation. Experimental results demonstrate that our SH2, requiring no additional data or models, can effectively help LLMs elicit factual knowledge and distinguish hallucinated contexts. Significant and consistent improvements are achieved by SH2 for LLaMA-7b, LLaMA2-7b and Mistral-7b on multiple hallucination tasks.  [^1] We release our code at https://github.com/0-KaiKai-0/SH2.
@@ 1 Introduction
Depending on massive training corpora, large language models (LLMs) have made tremendous progress in natural language understanding and text generation (Touvron et al., 2023a, Jiang et al., 2023, OpenAI, 2023). However, during reasoning and generation, LLMs could suffer from hallucinations and generate non-factual answers (Zhang et al., 2023b).
@@ 2 Related Work
The idea of using extra decoding steps while predicting harder tokens can be dated back to Adaptive Computation Time proposed by Graves (2017). Shin et al. (2020) introduce AUTOPROMPT to combine the original inputs with trigger tokens to elicit knowledge from pretrained models. Wiegreffe et al. (2022) investigate that additional steps to generate rationales could lead to a more faithful model. More recently, Goyal et al. (2023) have also demonstrated that for LLMs, inserting extra computation steps to allow the model to pause before generation can enhance the performance on question answering and reasoning tasks. Nevertheless, their method only works when the model is both pre-trained and finetuned with pauses.
@@ 3.2 Relation between Key Tokens and Factual Information
It can be observed from Figure 2 that content words such as adjectives (JJ), nouns (NN), proper nouns (NNP), adverbs (RB) and conjugated verbs (VBD, VBG, VBP) are more frequent in the hardest words we extract, as the light color in the heatmap indicates. On the other hand, function words such as conjunctions (CC), determiners (DT), and prepositions (IN) are less frequent, as indicated by the darker colors. [^2] The base form of the verbs (VB) seems to be an exception. It is probably because VBs have less information load than conjugated verb forms, as the latter also encode tense, aspect, and other information about the verb. This suggests that the normalized top-$\eta$ recall is positively correlated with factual informativeness. Therefore, picking out key tokens that carry more content, rather than those that serve as grammatical functions, could intuitively help the model to focus on factual information. Our method can thus be understood as a way to “highlight” the content words to LLMs, potentially making them less likely to hallucinate, the effect of which will be evaluated in our experiments. More cases will be given in Appendix D.2.
@@ 3.3 Construction of Hesitations
First, construct a candidate key-token set by selecting $\eta\cdot n$ tokens with the lowest prediction probabilities of $\hat{P}(X)$, where $\eta\in(0,1)$ is the sampling proportion we preset for $X$. Then, derive the target key-token set $T(X)$ by randomly drop out $\lambda\cdot\eta\cdot n$ tokens from the candidate set, where $\lambda\in[0,1]$ is the drop-out rate.
For long documents, to avoid $T(X)$ got dominated by one or two POS tags, we introduce the drop-out rate $\lambda$ to randomly select key tokens among a larger pool of candidates, whose size is determined by $\eta$. As a result, $(1-\lambda)\cdot\eta\cdot n$ tokens are retained in $T(X)$ to fill as input during the hesitation steps.
@@ 4.2 Experimental Settings
We apply our SH2 on LLaMA-7b (Touvron et al., 2023a), LLaMA2-7b (Touvron et al., 2023b) and Mistral-7b Jiang et al. (2023). We compare it with other SOTA (state-of-the-art) methods, including Alpaca (Taori et al., 2023), ITI (Li et al., 2023b), 13b-CD (Chuang et al., 2023) and DoLa (Chuang et al., 2023). All of these baselines use LLaMA-7b as their backbone. It should be noted that ITI uses 5% samples of TruthfulQA in few-shot prompting. Additionally, LLaMA-13b is used in 13b-CD to be contrasted with the 7b model. We implement Alpaca and DoLa following the official instructions and report the evaluation results of our implementation.
@@
Since there are only about a dozen tokens in each question of TruthfulQA, the sample proportion $\eta$ of LLaMA-7b is set to be 10% and 40% for the discrimination track and the generation track respectively. And the drop-out rate $\lambda$ is set to be 0. While for HaluEval-Sum, which has about a thousand tokens in each document, $\eta$ and $\lambda$ are set to be 6% and 0.33 for LLaMA-7b. The settings of the hyper-parameters $\eta$, $\lambda$ and $\alpha$ are summarized in Appendix A.
@@ 4.3 Main Results
The results on the three benchmarks are shown in Table 1. Our SH2 obtained significant and consistent improvements for LLaMA-7b, LLaMA2-7b and Mistral-7b. When compared with other SFT or decoding reformulation methods, our method outperforms them in most of the metrics of the benchmarks. There is no additional data or model used in our SH2. We ask models to select the hardest tokens and hesitate on them. Although LLaMA2 and Mistral have been aligned to generate truthful answers during its training procedure, they still benefit greatly from our inference-time method.
@@ 4.4 Bias in HaluEval-Sum
The great gap between Acc-A and Acc-H denotes LLMs’ bias in hallucinated summaries and right summaries even for LLaMA2-7b. While our method effectively alleviates the problem.
@@
Nevertheless, our method successfully mitigates this imbalance and benefits the general performance. It can advance the discriminant ability of models comprehensively.
(a) Average scores (%) for different highlighted tokens with the effective sampling proportion $\eta^{\prime}$ ranging from 1% to 8%. The errorbar denotes standard deviations.
(b) Harmonic accuracy (%) for different highlighted tokens with respect to the effective sampling proportion $\eta^{\prime}$.
Figure 3: Different choices of highlighted tokens by LLaMA2-7b.
@@ 5.1 Choices of Highlighted tokens
In contrast to key tokens that are the hardest to predict, we sample the easiest tokens with the highest prediction probability. Additionally, we also randomly sample the same number of tokens for comparison. We conduct experiments with LLaMA2-7b on 1000 samples of HaluEval-Sum. The effective sampling proportion $\eta^{\prime}=(1-\lambda)\cdot\eta$ ranges from 1% to 8% with the step of %1. Average scores and standard deviations are calculated for the three choices of highlighted tokens.
Figure 4: Precision, recall and F1 scores (%) for different highlighted tokens with the effective sampling proportion $\eta^{\prime}$ ranging from 1% to 8%. The errorbar denotes standard deviations. The scores of the vanilla LLaMA2-7b are obtained by evaluating on the whole dataset of HaluEval-Sum.
Moreover, we compare precision, recall and F1 scores for the three choices of highlighted tokens in Figure 4. The disturbance brought by highlighting the easiest tokens or random tokens hurts the precision of LLMs, while the hardest tokens are beneficial. It can also be concluded from the figure that the manner of hesitations helps mitigate the bias in HaluEval-Sum.
5.2 Manners of Hesitations
In addition to underlining key tokens in the input text, we can also repeat the text or pause. Experiments are conducted on TruthfulQA with LLaMA-7b. For the repetition manner, we can repeat the questions in hesitations. While for the pausing manner, key tokens $T_{2}(X)$ are replaced by several pause words (".") in hesitations. The performance of pausing hesitations is evaluated with 3, 6, 9 and 12 pause words. MC2 scores with different numbers of pause words and the adjustment hyper-parameter $\alpha$ are shown in Figure 6. 6 pause words are used in hesitations in the following study.
@@ 5.3 Effect of Contrastive Decoding
Figure 5: Effect of contrastive decoding on hesitations. The dashed line in gray represents the MC2 score of the vanilla LLaMA-7b. Dashed lines in other colors represent the MC2 scores for the standard contrastive decoding with different numbers of highlighted tokens.
Our method does not use the original form of contrastive decoding. The difference of confidence brought by hesitations, which is the item of contrastive decoding, is traded off with the base probability. We conduct experiments on TruthfulQA with different values of the adjustment hyper-parameter $\alpha$ to study the effect of contrastive decoding (CD).
The effect of contrastive decoding with different numbers of highlighted tokens is shown in Figure 5. Since there are only about a dozen tokens in each question of TruthfulQA, the hesitation contains only one token when sampling the top 10% hardest tokens. LLaMA-7b benefits a lot from the one or two highlighted tokens.
When $\alpha$ is set to be 0, the model directly decodes with the input text followed by the hesitation. The MC2 score even drops a little compared to the vanilla LLaMA-7b. As more weight is put on the difference brought by hesitations, the model obtains a more significant improvement. Specifically, when the weight is big enough to ignore the base probability, the decoding procedure is equivalent to standard contrastive decoding without the base probability as the equation (5). It can be concluded from the figure that, the highlighted tokens make a difference in the confidence to identify factual knowledge.
@@ 6 Conclusion
In this paper, we study the bias in capturing important factual information for large language models and introduce a novel inference-time method SH2. Our method proposes to highlight the key tokens that are hard for LLMs to predict and construct hesitations with these informative tokens. To help LLMs capture critical semantic information, the decoding procedure is reformulated with the probability difference brought by hesitations. Comprehensive experiments and analysis are conducted on multiple tasks. SH2 significantly improves the truthfulness of LLMs with no need for external data or models. In the age of LLMs, our SH2 could inspire a new paradigm for faithful LLMs.
Limitations
Existing research on LLMs’ hallucination does not pay much attention to other dimensions of the generation quality. Despite experiments on the generation track of TruthfulQA, we have yet to explore the diversity and soundness of the generated contents by our SH2, which is closely related to the generalization ability. It is worth studying how to optimize LLM’s truthfulness and generalization simultaneously.
Besides, our method is an inference-time method without leveraging external data or models. Consequently, it could be promising to integrate our method with other data-enhanced or model-enhanced methods. The idea of constructing hesitations can also be applied in retrieval augmentation approaches.
Ethics Statements
Our work pertains to large language models’ hallucinations. In this work, we use only publicly available data and artifacts. There are no ethical issues in our paper, including its motivation and experiments.
Appendix A Hyper-parameter Settings
The settings of the hyper-parameters $\eta$, $\lambda$ and $\alpha$ are summarized in Table 4. The drop-out rate $\lambda$ is set to be 0 for the discrimination track and the generation track of TruthfulQA, whose questions are of a dozen tokens. It is set to be 0.33 for FACTOR and HaluEval-Sum, because these benchmarks have longer documents.
@@ Appendix B Multiple-Choice Metrics
Appendix C Constrastive Decoding
C.1 Effect of CD with Pause Words
Figure 6: Effect of contrastive decoding on hesitations. The dashed line in gray represents the MC2 score of the vanilla LLaMA-7b. Dashed lines in other colors represent the MC2 scores for the standard contrastive decoding with different numbers of pause words.
As shown in Figure 6, we also studied the effect of contrastive decoding with different numbers of pause words. A similar conclusion can be observed from the figure.
C.2 Case Study
A case study is given in Figure 7 to illustrate the effect of our self-highlighted hesitations and contrastive decoding. We calculate the probability given by LLaMA-7b. $p_{\theta}(Y|X)$ and $p_{\theta}(Y|X^{\prime})$ stand for standard decoding without and with the hesitation for each option. The difference in probability are derived by $\frac{p_{\theta}(Y|X^{\prime})}{p_{\theta}(Y|X)}$.
It can be learned from the figure that, with hesitations, LLaMA-7b is prone to assign higher probabilities to correct answers and lower probabilities to false answers. The model is more confident in selecting correct answers and rejecting false answers. It demonstrates that CD helps LLMs distinguish correct answers from false ones. It plays a role in separating the positive and negative answers.
Appendix D Key Tokens
D.1 Normalized Top-$\eta$ Recall
We also calculated the normalized top-$\eta$ recall between the hardest part and the whole document for the top 20 most frequent POS tags by LLaMA-7b and Mistral-7b through the equation (3). The heatmaps are shown in Figure 8. The same statistical laws can be concluded as those of LLaMA2-7b in Section 3.2.
(a) The normalized top-$\eta$ recall of LLaMA-7b.
(b) The normalized top-$\eta$ recall of Mistral-7b.
Figure 8: The normalized top-$\eta$ recall for different POS tags by LLaMA-7b and Mistral-7b.
D.2 Visualization of Token Probabilities
@@
(c) Visualization of token probabilities estimated by Mistral-7b.
Figure 9: Visualization of token probabilities estimated by LLaMA-7b, LLaMA2-7b and Mistral-7b on case 1. The color coding in the figure represents the token generation probabilities. Tokens with lower probabilities are colored more red, while tokens with higher probabilities are colored more green.
@@
(c) Visualization of token probabilities estimated by Mistral-7b.
Figure 10: Visualization of token probabilities estimated by LLaMA-7b, LLaMA2-7b and Mistral-7b on case 2. The color coding in the figure represents the token generation probabilities. A spectrum from red to green is utilized to represent a range from lower to higher generation probabilities, respectively.
@@
(c) Visualization of token probabilities estimated by Mistral-7b.
Figure 11: Visualization of token probabilities estimated by LLaMA-7b, LLaMA2-7b and Mistral-7b on case 3. The color coding in the figure represents the token generation probabilities. Tokens with lower probabilities are colored more red, while tokens with higher probabilities are colored more green.
Figure 9 to Figure 11 illustrate three text examples where the background color of each token denotes its generation probability given by LLaMA-7b, LLaMA2-7b, and Mistral-7b. A spectrum from red to green is utilized to represent probabilities ranging from low to high.
These examples reveal a tendency for lower probability tokens to consist of content words predominantly. However, instances exist where determiners and prepositions also exhibit lower probabilities, as exemplified in Figure 9a with the words “all” and “on” in line 5, and the word “in” in line 14. This can be attributed to the interchangeable use of different words to convey identical semantic content at the beginning of a sentence or sub-clause.
Besides, tokens with higher probabilities are typically from function words. However, exceptions are noted with certain content words demonstrating high probabilities due to the extensive knowledge memorized by LLMs, such as “typhus” and “15” in the first line of Figure 11a. Additionally, some non-initial tokens, such as “nesday” and “CC” in lines 1 and 3 of Figure 10a, exhibit high generation probabilities. It reflects the determinative role of initial tokens in setting the context for subsequent non-initial tokens.
Comparing the prediction probabilities of LLaMA-7b, LLaMA2-7b, and Mistral-7b, we did not observe significant difference. Although LLaMA2-7b and Mistral-7b demonstrate superior performance in various benchmarks compared to LLaMA-7b, there is still considerable room for improvement in their grasp of factual knowledge. Therefore, it is beneficial to incorporate hesitations to make the model pay more attention to these informative key tokens.